simplify: 15-5i2+3i32
Answer by stanbon(75887)

simplify: 15-5i^2+3i^32
---
i^2 = -1
i^32 = (i^4)^8 = 1*8 = 1
---
Your Problem = 15-5*-1+3*1 = 15+5+3 = 23
===========================
Cheers,
Stan H.
==============